(4 points) Consider the error in using the approximation sin:) on the interval (-1,1]. (a) Reasoning informally, on what inteI just need help with C and D. I tried (0,1) and (-1,0) for C and 1/6! for D
Answer #1

(c) overestimated [-1,0) and underestimated from (0,1] since sin1=0.84 while x-x^3 at x=1 is 0.8333 so underestimated in (0,1] similarly sin(-1)=-0.84 while x-x^3 at x=-1 is -0.8333 so 0verestimated [-1,0).

(d) smallest error bound= 1/5!

Since sinx=x+0x^2+(x^3)/6 +0x^4 . So R4(x)=1/5!